A 125 m long transmission wire is stretched between two vertical posts. It carries a current of 35.0 A and experiences a force of 0.25 N due to Earth's magnetic field. What is the value of Earth's magnetic field at the location of the wire if the wire is 60° with it?

1
Expert's answer
2022-03-22T11:00:13-0400

The magnetic field force is given by

"F=BIl\sin\theta"

Hence, the strength of magnetic field

"B=\frac{F}{Il\sin\theta}\\
=\frac{0.25}{35.0*125*\sin60^\circ}=6.6*10^{-5}\:\rm T"